OpenStudy (anonymous):

Choose the equation below that represents the line passing through the point (2, -4) with a slope of 1/2

OpenStudy (allank):

Hint: The equation of a line in point-slope form is y=m(x-x1)+y1 where (x1,y1) is a point on the line and m is the gradient of the line.

OpenStudy (allank):

Can you calculate the equation of the line now? @smantha

OpenStudy (anonymous):

no i still dont understand

OpenStudy (allank):

This is the equation of a line in point-slope form: y=m(x-x1)+y1 . There are other forms of the equation of a line, but point-slope will work best. In this question, the gradient (m)=1/2. We have also been told that the line passes through the point (2,-4) thus x1=2 and y1=-4 . We now fix in the values of m, x1 and y1 into the equation y=m(x-x1)+y1 y=1/2(x-2)-4 Simplify that and you will get the equation of the line.

OpenStudy (allank):

Does that make more sense?

OpenStudy (anonymous):

yes thank you

OpenStudy (allank):

You're welcome. :)